# Exploring the Low-Pass Filtering Behavior in Image Super-Resolution

### v1

@@ Abstract
Deep neural networks for image super-resolution (ISR) have shown significant advantages over traditional approaches like the interpolation. However, they are often criticized as ‘black boxes’ compared to traditional approaches with solid mathematical foundations. In this paper, we attempt to interpret the behavior of deep neural networks in ISR using theories from the field of signal processing. First, we report an intriguing phenomenon, referred to as ‘the sinc phenomenon.’ It occurs when an impulse input is fed to a neural network. Then, building on this observation, we propose a method named Hybrid Response Analysis (HyRA) to analyze the behavior of neural networks in ISR tasks. Specifically, HyRA decomposes a neural network into a parallel connection of a linear system and a non-linear system and demonstrate that the linear system functions as a low-pass filter while the non-linear system injects high-frequency information. Finally, to quantify the injected high-frequency information, we introduce a metric for image-to-image tasks called Frequency Spectrum Distribution Similarity (FSDS). FSDS reflects the distribution similarity of different frequency components and can capture nuances that traditional metrics may overlook. Code for this paper can be found in: https://github.com/RisingEntropy/LPFInISR .
@@ 1 Introduction
Figure 1: $I$ is an image in which only the central pixel is 1 and the other pixels are 0. What will the result be like if $I$ is super-resolved using a neural network, A, B, C or D? Surprisingly, the answer is A. We name this phenomenon as the sinc phenomenon. In this paper, we give a possible explanation for this phenomenon.
The goal of image super-resolution (ISR) is to reconstruct low-resolution (LR) images into high-resolution (HR) images through various techniques. In recent years, with advances in deep learning, growing ISR methods using neural networks are proposed, bringing the development of ISR into a new level. While impressive results persistently arise, the theoretical underpinnings of ISR in neural networks remain largely unexplored, leading to criticism that they are considered black boxes. In comparison, traditional methods, such as interpolation or filtering, have strong interpretability. Despite the principles of traditional methods and neural networks are different, we can still attempt to explain the behavior of neural networks using the theories of traditional methods. In this paper, following this line of thought, we successfully utilize theories from traditional signal processing techniques to explain the performance of neural networks in ISR task.
The essence of the ISR task lies in the process of upsampling a two-dimensional signal. In traditional signal processing theory [27, 17], a feasible method for upsampling involves restoring a discrete low-sampling-rate signal to a continuous signal using a low-pass filter, and then sampling the continuous signal at a higher rate to obtain a high-sampling-rate signal. An intriguing aspect of this process is that when we try to upsampling a Dirac $\delta$ signal, we will finally get a sinc signal since the sinc signal is the time-domain waveform of a low-pass filter, (for details about this, please refer to Sec. 3 ). Given this, we can conjecture: if neural networks exhibit similar behavior, then when attempting super-resolution on a Dirac $\delta$ signal, the resultant outcome would also be a sinc signal. As shown in Fig. 1, we indeed observe this phenomenon, and we name it as ‘the sinc phenomenon’. This phenomenon establishes a connection between traditional signal processing theory and the interpretability of neural networks, thus helping us form a deeper understanding of neural networks.
@@
Given that the non-linear component is injecting high-frequency information, there is a pressing need for a metric to quantitatively describe the extent of the injected high frequencies. Previous metrics, like PSNR, SSIM [25] and LPIPS [36], have not approached the evaluation of images from a frequency perspective. Therefore, we propose the frequency spectrum distribution similarity (FSDS), a metric that evaluate image quality based on the power distribution in the frequency spectrum.
@@
  • We report an intriguing phenomenon: the impulse response of image super-resolution (ISR) neural network is a sinc function, representing the temporal waveform of a low-pass filter. We name it after ‘the sinc phenomenon’. This observation aids in establishing a connection between signal processing theory and neural networks.
  • In order to further explain the performance of neural networks in ISR tasks through this phenomenon, we introduce HyRA. HyRA considers the neural network as a parallel combination of a linear system and a non-linear system with a zero impulse response. It points out that the linear system operates as a low-pass filter, while the non-linear system injects high-frequency information.
  • To quantitatively describe the injection of high frequencies, we introduce the FSDS metric. FSDS can reflect high-frequency distortions that previous metrics fail to capture.
@@ 2.1 Super Resolution Using Neural Networks
Recent review articles in ISR include fixed-scale super-resolution [33] and arbitrary-scale super-resolution review [14]. There are many architectures of mainstream ISR backbone networks, including CNN-style backbones [13, 39, 9, 38, 2], transformer-style backbones [12, 24] and GAN-style backbone networks [23], etc. Depending on these backbones, researchers have proposed many modules with various functions. For example, ArbSR [22] can expand a fixed-scale super-resolution network to an arbitrary-scale ISR network, LTE [10] can enhance local textures, etc. What worth mentioning is that LIIF [5] introduces implicit neural representation into ISR for the first time, bringing a new approach for ISR. This paper mainly focuses on approaches that utilize CNN-style or transformer-style backbones. Except for network architectures, numerous datasets have been proposed to facilitate further research. Commonly used datasets for ISR encompass Set5 [3], Urban100 [8], Flickr2K [34], SCI1K [32], DIV2K [1], etc. We evaluate the effectiveness of our proposed FSDS metric on DIV2K dataset. The large size of the DIV2K dataset contributes to increased reliability in our conclusions.
@@ 2.2 Explaining the Behavior of Neural Networks
Despite neural networks are often criticized as ‘black boxes,’ predecessors have made remarkable efforts to mitigate this situation. Various previous researches have proposed many methods to analyze the behavior of neural networks. Since Sundararajan et al. [20] introduce the integrated gradients (IG) for attribution in classification tasks, numerous researchers have expanded this method to various domains, broadening the scope of attribution beyond classification tasks. Based on IG, Gu & Dong [6] propose LAM to analyze the impact of the local patch on the entire ISR outcome. However, such a method requires manually determined hyper-parameters and baselines, thus introducing subjectivity. Several notable analysis methods leveraging Fourier transform have been explored in literature [31, 29, 30, 40]. Notably, Xu et al. [31] advocate for the Frequency-Principle, asserting its relevance to both convolutional neural networks (CNNs) and fully-connected deep neural networks. According to their proposition, these networks inherently adhere to the Frequency-Principle, wherein training data is systematically acquired in a sequential manner, progressing from low to high frequency. In a departure from conventional approaches, the HyRA method distinguishes itself by employing impulse response to probe the potential mechanisms of deep neural networks in the context of ISR tasks.
@@ 3 Preliminaries
Section B.1-Section B.3 provide a brief overview for readers who are not familiar with the signals and systems. Section B.1 introduces the concepts of signals and systems, along with the computation of responses in Linear Time-Invariant (LTI) systems. Section B.2 covers the processes of signal sampling and reconstruction. Section B.3 delves into the phenomenon of spectrum aliasing, a factor contributing to the ill-posed nature of the ISR task. And we will introduce applying low-pass filter for ISR in this section.
We can employ signal recovery methods to achieve image super-resolution (ISR). Initially, we conceptualize a discrete image as a series of impulse trains in a two-dimensional continuous space, with varying densities representing different resolutions. Then, for the low-resolution image, we begin by implementing low-pass filtering, following the procedure outlined in Section B.2, to obtain the continuous image $I^{\mathrm{cont}}$, This process can be mathematically described as:
@@
where * denotes convolution, $I^{LR}_{x,y}$ is the low resolution image with variant $x,y$ and $I^{\mathrm{cont}}$ is the continuous signal. $sinc^{\omega}_{x,y}$ is a two-dimensional sinc function with parameter $\omega$, namely it’s frequency spectrum is an ideal low-pass filter with a passband of $0\sim\omega$. Subsequently, we sample the ‘conceptually continuous signal’ at an elevated sampling rate to acquire a more densely populated two-dimensional sequence of impulse trains, that is an image with higher resolution denoted as $I^{SR}$:
@@
In the equation, $s^{\Delta X,\Delta Y}_{x,y}$ denotes the two-dimensional impulse train with intervals of $\Delta X$ in $x$ axis and $\Delta Y$ in $y$ axis.
@@ 4.1 Hybrid Response Analysis (HyRA)
In this section, we describe the proposed Hybrid Response Analysis (HyRA), which treating the neural network as a combination of a linear system and a non-linear system. Through the impulse response, we can calculate a linear time invariant (LTI) system’s output from any input using the convolution operation (see Section B.1). However, since neural networks are nonlinear systems, we cannot apply convolution to analyze them. To further study the network features, we need to split it into a linear system and a non-linear system, i.e., HyRA. The core concept HyRA is illustrated in Fig. 3. We denote an ISR network as $N(I)$, where $I$ is the input image. $N(I)$ is a non-linear system that can be expressed as the sum of a linear system and a non-linear system:
@@
This is the invariance of a single convolution layer. ∎
Through experiments (please see Fig. 15), we demonstrate that the linear system of ISR neural networks, when analyzed using the HyRA method, indeed behaves as a ‘spatially invariant system.’
@@
where * means the convolution operation and $H(\delta)$ is the impulse response. Although the response of the non-linear component cannot be directly computed, if we obtain the final output of the neural network, the non-linear part can be deduced by subtracting the response of the linear component from the final output, namely the non-linear response can be computed as:
@@ 4.1.1 $H(I)$ is a low-pass filter
Figure 4: Top row: one ISR image by EDSR [13] can be viewed as the summation of a linear response obtained by convolving impulse response with the input and the non-linear response gained by subtracting linear-part from ISR result. Second row: the corresponding frequency spectrum amplitude of the top row. Third row: the corresponding frequency spectrum phase of the top row. It is clear that there exists significant phase compensation.
In Sec. 3, we mention that a simple low-pass filter achieves ISR functionality. Do neural networks possess low-pass filters internally? If this hypothesis is valid, according to the principle of HyRA, when we input a Dirac $\delta$ signal into the neural network $N(I)$, the output should be the impulse response of the low-pass filter, namely the sinc function (please recall Section B.1 and Tab. 3). In the experiment section (Section 5.2), we conduct tests on three mainstream ISR backbones and some derived methods. We find that their impulse responses are sinc functions [^2] Strictly speaking, it is a windowed sinc function. Regarding the windowing operation, please refer to Appx. E. Now, with both the impulse response and spatial invariance property, we can compute the response of the linear system $H(I)$ to any input through convolution:
@@ 4.1.2 $G(I)$ Injects high-frequency information
As mentioned in Section 4.1.1, the non-linear component also plays a crucial role in compensating for the distortion introduced by $H(I)$. Examining the response of $G(I)$, we note that it also exhibits grid-like distortions, matching those in the response of $H(I)$. This allows for the cancellation of the grid-like distortions, achieving the final goal of ISR. As shown in Fig. 4, upon observing the frequency spectrum, bright spots corresponding to the amplitude spectrum of $H(I)$ exist in all four corners of the amplitude spectrum of $G(I)$. However, the phase spectrum of $G(I)$ is in compensation of the phase spectrum of $H(I)$, indicating such
@@ 4.2 Frequency Spectrum Distribution Similarity (FSDS)
In this section, we introduce the FSDS metric to quantitatively describe the extent of the ‘injected high frequencies’ as discussed in Section 4.1.2.
@@ 4.2.1 Motivation and Method
Additionally, we’ve noted that the frequency domain distribution in the ISR field can significantly impact downstream applications [35, 28]. Consequently, we propose that evaluating the ISR effectiveness of a network requires a thorough assessment of its performance in the frequency spectrum. This involves examining the similarity in frequency spectra between the low-resolution image and the high-resolution image. The Frequency Spectrum Distribution Similarity (FSDS) metric integrates the power distribution maps of the spectra for both images. The difference is then calculated to generate an error map, and the total sum of its absolute values is computed.
@@
where $E(I^{\mathrm{HR}})$ and $\sigma(I^{\mathrm{HR}})$ are the mean and variance of $I^{\mathrm{HR}}$ respectively. Similarly, we perform a Fourier transform on the ISR image $I^{\mathrm{SR}}_{j\omega_{1},j\omega_{2}}$. It is worth noting that unlike other metrics, such as PSNR and SSIM[25], which do not incorporate normalization, FSDS is specifically designed to accentuate numerical variations due to its emphasis on numerical changes rather than absolute numerical values. Then, the complex integration of the two spectra is performed, providing the power distribution map $D^{\mathrm{HR}}$, which is defined as:
@@
Similarly, we can obtain $D^{SR}$. Subsequently, the difference between $D^{HR}$ and $D^{SR}$ is calculated, resulting in the difference map $D^{\mathrm{diff}}$ of their power distribution:
@@
where ||$\cdot$|| ⋅ | represents taking the magnitude of a complex number. Considering a more concise description of a larger dynamic range, logarithm is taken. A larger FSDS value indicates that the two images are closer, thereby suggesting better ISR results. It is essential to note that the rationale behind FSDS differs from commonly used ISR metrics, such as PSNR and SSIM.

### v4

@@ Abstract
Deep neural networks for image super-resolution (ISR) have shown significant advantages over traditional approaches like the interpolation. However, they are often criticized as ‘black boxes’ compared to traditional approaches with solid mathematical foundations. In this paper, we attempt to interpret the behavior of deep neural networks in ISR using theories from the field of signal processing. First, we report an intriguing phenomenon, referred to as ‘the sinc phenomenon.’ It occurs when an impulse input is fed to a neural network. Then, building on this observation, we propose a method named Hybrid Response Analysis (HyRA) to analyze the behavior of neural networks in ISR tasks. Specifically, HyRA decomposes a neural network into a parallel connection of a linear system and a non-linear system and demonstrates that the linear system functions as a low-pass filter while the non-linear system injects high-frequency information. Finally, to quantify the injected high-frequency information, we introduce a metric for image-to-image tasks called Frequency Spectrum Distribution Similarity (FSDS). FSDS reflects the distribution similarity of different frequency components and can capture nuances that traditional metrics may overlook. Code, videos and raw experimental results for this paper can be found in: https://github.com/RisingEntropy/LPFInISR.
@@ 1 Introduction
Figure 1: $I$ is an image in which only the central pixel is 1 and the other pixels are 0. What would the result look like if image I is super-resolved using a neural network, A, B, C, or D? Surprisingly, the answer is A. We name this phenomenon as the sinc phenomenon. In this paper, we give a possible explanation for this phenomenon.
The goal of image super-resolution (ISR) is to reconstruct low-resolution (LR) images into high-resolution (HR) images through various techniques. In recent years, with advances in deep learning, growing ISR methods using neural networks are proposed, bringing the development of ISR into a new level. While impressive results persistently arise, the mechanism under ISR networks remain largely unexplored, leading to criticism that they are considered black boxes. In comparison, traditional methods, such as interpolation or filtering, have strong interpretability. Despite the principles of traditional methods and neural networks are different, we can still attempt to explain the behavior of ISR networks using theories from traditional methods. In this paper, following this line of thought, we successfully utilize theories from the field of signal processing techniques to explain the performance of neural networks in the ISR task.
The target of the ISR task is to upsample a two-dimensional signal. In traditional signal processing theory [17, 18], a feasible method for upsample involves restoring a discrete low-sampling-rate signal to a continuous signal using a low-pass filter, and then sampling the continuous signal at a higher rate to obtain a high-sampling-rate signal. An intriguing aspect of this process is that when we try to upsample a Dirac $\delta$ signal, we will finally get a sinc signal since the sinc signal is the time-domain waveform of a low-pass filter, (for details about this, please refer to Sec. 3 ). Given this, we can conjecture: if neural networks exhibit similar behavior, then when attempt to super-resolve a Dirac $\delta$ signal, the resultant outcome would also be a sinc signal. As shown in Fig. 1, we indeed observe this phenomenon, and we name it as ‘the sinc phenomenon’. This phenomenon establishes a connection between traditional signal processing theory and the interpretability of neural networks, thus helping us form a deeper understanding of ISR networks.
@@
Given that the non-linear component is injecting high-frequency information, there is a pressing need for a metric to quantitatively describe the extent of the injected high frequencies. Previous metrics, like PSNR, SSIM [26] and LPIPS [36], have not approached the evaluation of images from a frequency perspective. Therefore, we propose the frequency spectrum distribution similarity (FSDS), a metric that evaluates image quality based on the power distribution in the frequency spectrum.
@@
  • We report an intriguing phenomenon: the impulse responses of image super-resolution (ISR) networks are sinc functions, representing the temporal waveform of a low-pass filter. We name it the ’sinc phenomenon’. This observation helps to establish a connection between signal processing theory and neural networks. Moreover, we find that for a network, the more similar the impulse response is to the sinc function, the better performance it produces.
  • In order to further explain the performance of neural networks in the ISR task through this phenomenon, we introduce HyRA. HyRA considers the neural network as a parallel combination of a linear system and a non-linear system with a zero impulse response. It points out that the linear system operates as a low-pass filter, while the non-linear system injects high-frequency information.
  • To quantitatively describe the injection of high frequencies, we introduce the FSDS metric. FSDS measures image quality using frequency spectrum produced by FFT and can reflect high-frequency distortions that previous metrics fail to capture.
@@ 2.1 Super Resolution Using Neural Networks
Recent review articles in ISR include fixed-scale super-resolution [33] and arbitrary-scale super-resolution review [14]. There are various architectures of mainstream ISR backbone networks, including CNN-style backbones [13, 38, 9, 37, 2], transformer-style backbones [12, 25] and GAN-style backbone networks [24], etc. Based on these backbones, researchers have proposed quantitative modules with various functions. For example, ArbSR [23] can expand a fixed-scale super-resolution network to an arbitrary-scale ISR network, LTE [10] can enhance local textures, etc. What worth mentioning is that LIIF [5] introduces implicit neural representation into ISR for the first time, bringing a new approach for ISR. This paper mainly focuses on approaches that utilize CNN-style or transformer-style backbones. Except for network architectures, numerous datasets have been proposed to facilitate further research. Commonly used datasets for ISR includes Set5 [3], Urban100 [8], Flickr2K [34], SCI1K [32], DIV2K [1], etc. We evaluate the effectiveness of our proposed FSDS metric on DIV2K dataset. The large size of the DIV2K dataset contributes to increased reliability in our conclusions.
@@ 2.2 Explaining the Behavior of Neural Networks
Despite neural networks are often criticized as ‘black boxes,’ predecessors have made remarkable efforts to mitigate this situation. Various previous researches have proposed plenty of methods to analyze the behavior of neural networks. Since Sundararajan et al. [21] introduce the integrated gradients (IG) for attribution in classification tasks, numerous researchers have expanded this method to various domains, broadening the scope of attribution beyond classification tasks. Based on IG, Gu & Dong [6] propose LAM to analyze the impact of the local patch on the entire ISR outcome. However, such a method requires manually determined hyper-parameters and baselines, thus introducing subjectivity. Several notable analysis methods utilizing the Fourier transform have been explored in the literature [31, 29, 30, 39]. Notably, Xu et al. [31] propose the Frequency-Principle, claiming its relevance to both convolutional neural networks (CNNs) and fully-connected deep neural networks. According to their proposition, these networks inherently adhere to the Frequency-Principle, wherein training data is systematically acquired in a sequential manner, progressing from low to high frequency. Unlike previous approaches these approaches, HyRA distinguishes itself by employing impulse response to probe the potential mechanisms of deep neural networks in the context of the ISR task.
@@ 3 Preliminaries
Section B.1-Section B.3 provide a brief overview of signal processing concepts for readers who are not familiar with it. Section B.1 introduces the concepts of signals and systems, along with the computation of responses in Linear Time-Invariant (LTI) systems. Section B.2 covers the processes of signal sampling and reconstruction. Section B.3 delves into the phenomenon of spectrum aliasing, a factor contributing to the ill-posed nature of the ISR task. And we will introduce applying low-pass filter for ISR here.
We can employ signal recovery methods to achieve image super-resolution (ISR). Initially, we conceptualize an image as a series of impulse trains in a two-dimensional continuous space, with varying densities representing different resolutions. Then, for the low-resolution image, we begin by implementing low-pass filtering, following the procedure outlined in Section B.2, to obtain the continuous image $I^{\mathrm{cont}}$, This process can be mathematically described as:
@@
where * denotes convolution, $I^{LR}_{x,y}$ is the low resolution image with variant $x,y$ and $I^{\mathrm{cont}}_{x,y}$ is the continuous signal. $sinc^{\omega}_{x,y}$ is a two-dimensional sinc function with parameter $\omega$ [^2] Please refer to Tab. 3 for Fourier transform pairs, whose frequency spectrum is an ideal low-pass filter with a passband of $0\sim\omega$. Subsequently, we sample the ‘conceptually continuous signal’ at an elevated sampling rate to acquire a more densely populated two-dimensional sequence of impulse trains, i.e., an image with higher resolution denoted as $I^{SR}$:
@@
In the equation, $s^{\Delta X,\Delta Y}_{x,y}$ denotes the two-dimensional impulse trains with intervals of $\Delta X$ in $x$ axis and $\Delta Y$ in $y$ axis.
@@ 4.1 Hybrid Response Analysis (HyRA)
In this section, we describe the proposed Hybrid Response Analysis (HyRA), which treats the neural network as a combination of a linear system and a non-linear system. Through the impulse response, we can calculate a linear time invariant (LTI) system’s output from any input using the convolution operation (see Section B.1). However, since neural networks are nonlinear systems, we cannot apply convolution to analyze them. To further explore the network features, we need to split it into a linear system and a non-linear system, i.e., HyRA. The core concept HyRA is illustrated in Fig. 3. We denote an ISR network as $N(I)$, where $I$ is the input image. $N(I)$ is a non-linear system that can be expressed as the sum of a linear system and a non-linear system:
@@
This is the invariance of a single convolution layer, and still holds for more layers. ∎
@@
where * means the convolution operation. Although the response of the non-linear component cannot be directly computed, if we obtain the final output of the neural network, the non-linear part can be deduced by subtracting the response of the linear component from the final output, namely the non-linear response can be computed as:
@@ 4.1.1 $H(I)$ is a low-pass filter
Figure 4: Top row: a super-resolved image by [13] can be viewed as the summation of a linear response obtained by convolving impulse response with the input and the non-linear response gained by subtracting linear-part from the ISR result. Second row: the corresponding frequency spectrum amplitude of the top row. Third row: the corresponding frequency spectrum phase of the top row. The phase compensation indicates that the non-linear part is compensating distortion.
In Sec. 3, we mention that a simple low-pass filter achieves ISR functionality. Do neural networks possess low-pass filters internally? If this hypothesis is valid, according to the principle of HyRA, when we input a Dirac $\delta$ signal into the neural network $N(I)$, the output should be the impulse response of the low-pass filter, i.e., the sinc function (please recall Section B.1 and Tab. 3). In the experiment section (Section 5.2), we conduct tests on three mainstream ISR backbones and some derived methods. We find that their impulse responses are sinc functions [^3] Strictly speaking, it is a windowed sinc function. Regarding the windowing operation, please refer to Appx. E. Now, with both the impulse response and spatial invariance property, we can compute the response of the linear system $H(I)$ to any input through convolution:
@@ 4.1.2 $G(I)$ Injects high-frequency information
As mentioned in Section 4.1.1, the non-linear component also plays a crucial role in compensating for the distortion introduced by $H(I)$. Examining the response of $G(I)$, we note that it also exhibits grid-like distortions, matching those in the response of $H(I)$. This allows for the cancellation of the grid-like distortions, achieving the final goal of ISR. As shown in Fig. 4, upon observing the frequency spectrum, bright spots corresponding to the amplitude spectrum of $H(I)$ exist in all four corners of the amplitude spectrum of $G(I)$. However, the phase spectrum of $G(I)$ is in compensation of the phase spectrum of $H(I)$, indicating that the grid-like distortion is ‘erased’ here.
@@ 4.2 Frequency Spectrum Distribution Similarity (FSDS)
In this section, we introduce the FSDS metric to quantitatively describe the so called the ‘injected high frequencies’ as discussed in Section 4.1.2.
4.2.1 Motivation and Method
@@
Additionally, we’ve noted that the frequency domain distribution in the ISR field can significantly impact downstream applications [35, 28]. Consequently, we propose that evaluating the ISR effectiveness of a network requires a thorough assessment of its performance in the frequency spectrum. This involves examining the similarity in frequency spectrum between the low-resolution image and the high-resolution image. The Frequency Spectrum Distribution Similarity (FSDS) metric integrates the power distribution maps of the spectrum for both images. The difference is then calculated to generate an error map, and the total sum of its absolute values is computed.
@@
where $E(I^{\mathrm{HR}})$ and $\sigma(I^{\mathrm{HR}})$ are the mean value and variance of $I^{\mathrm{HR}}$ respectively. Similarly, we perform a Fourier transform on the ISR image to obtain $I^{\mathrm{SR}}_{j\omega_{1},j\omega_{2}}$. It is worth noting that unlike other metrics, such as PSNR and SSIM[26], which do not incorporate normalization, FSDS is specifically designed to accentuate numerical variations due to its emphasis on numerical changes rather than absolute numerical values. Then, the complex integration of the two spectrum is performed, providing the power distribution map $D^{\mathrm{HR}}$, which is defined as:
@@
Similarly, we can obtain $D^{SR}$. Subsequently, the difference between $D^{HR}$ and $D^{SR}$ is calculated, providing a difference map $D^{\mathrm{diff}}$ of their power distribution:
@@
where ||$\cdot$|| ⋅ | represents taking the magnitude of a complex number. Considering a more concise description of a larger dynamic range, logarithm is taken. A larger FSDS value indicates that the two images are closer, thereby suggesting better ISR results.